If k(x) =6x^2-4x+2,find k(4+y) enter the expression in the simplest form. the terms of expression must be entered in descending order of degree k(4+y)=
Found 2 solutions by tjansen, stanbon:
Answer by tjansen(4)

Procedure:
Replace each x in the original equation with (4+y), then simplify.
k(4+y)=6(4+y)^2-4(4+y)+2 Replaced x with (4+y)
=6(16+8y+y^2)-16-4y+2 Squared (4+y) and distributed -4 over (4+y)
=96+48y+6y^2-14-4y Distributed 6 over (16+8y+y^2) and combined -16+2=-14
=6y^2+44y+82 END Combined like terms and put in descending order of degree.

Answer by stanbon(75887)

If k(x) =6x^2-4x+2,find k(4+y) enter the expression in the simplest form. the terms of expression must be entered in descending order of degree k(4+y)=
---
Just substitute 4+y for "x".
----
k(4+y) = 6(4+y)^2 - 4(4+y) + 2
======================================
Cheers,
Stan H.